OpenStudy (anonymous):

A square pyramid has a slant height of 17 inches and a height of 15 inches. What is its surface area if the square base has sides of length 16 inches?

OpenStudy (anonymous):

What do you mean by slant height?

OpenStudy (anonymous):

i have no idea,this is on my geometry homework

OpenStudy (anonymous):

736 is not an option though...

OpenStudy (anonymous):

I suggest going through your notes and figuring out what slant height is. Or searching on the internet somewhere. It's been my experience that trying to work out a maths problem when I don't know what the words mean is very hard!

OpenStudy (anonymous):

Can you give the options?

OpenStudy (anonymous):

What have you been learning so i know which method we are aiming for

OpenStudy (anonymous):

224 in2 480 in2 800 in2 864 in2

OpenStudy (anonymous):

this is lateral and surface area.

OpenStudy (anonymous):

The formula you are looking for is 2bs+b^2

OpenStudy (anonymous):

2x16x17+16x16

OpenStudy (anonymous):

=800

OpenStudy (anonymous):

b=base s=slant

OpenStudy (phi):

Here is the formula http://en.wikipedia.org/wiki/Pyramid_(geometry)#Surface_area

OpenStudy (anonymous):

Thank you guys.

OpenStudy (anonymous):

Phi that is a complex formula :P

OpenStudy (phi):

Find the area of the base (the bottom) which is a square 16*16 find the area of the sides. this is Perimeter of the base/2 times slant height Perimeter is 4 times the base side= 4*16= 64 64/2= 32 so 32*17 for the sides

OpenStudy (phi):

Take these things a little bit at a time.

OpenStudy (anonymous):

You must live a complicated life Phi :P

OpenStudy (phi):

No. Just take small steps. Can you find the area of the base (you know it is a square) with a side= 16

OpenStudy (phi):

Can you find the perimeter of the base? 4* side = 4*16

OpenStudy (anonymous):

Phi who you teaching?

OpenStudy (phi):

The formula says Perimeter/2 so 64/2 = 32 *who ever is curious

OpenStudy (anonymous):

That must be i then, continue please :P

OpenStudy (phi):

Area = Base area + P*S/2 where we already figured out Base area= 256 P is perimeter , we got 4*16= 64 P/2 =32 S is given as 17 so Area = 256 + 32*17 = 800

OpenStudy (phi):

It is no harder than following a recipe to cook a meal (actually easier, because you can't hurt yourself when dicing up the food)

OpenStudy (anonymous):

A question to your Highness, Isn't my formula less time consuming

OpenStudy (phi):

It looks fine. The important thing is (1) remember it (2) be able to use it.

OpenStudy (anonymous):

AyeAye Capt'n, Interesting techniques yours. Reminds me of someone cant remember who ::P

OpenStudy (anonymous):

Thanks for the intriguing method, it is easier to understand if your not into geometry